OpenStudy (anonymous):

use to substitute x-2y=3 4x-8y=12

OpenStudy (anonymous):

x-2y=3 x=3+2y do the same for the top 3+2y-2y=3 s0... 3=3 4x-8y=12 4(2y+3)-8y=12 8y+12-8y=12, so ...12=12